1. In view of the quantum number r dependence of the second correction factor,explain why the the real molecule s needed to describe the vibrational-rotational energy ofa
Answer #1

Step 1

A molecule can achieves both vibrational and rotational motions as well while atom can’t achieve it. The change in energy takes place when a molecule moves and movement of molecule takes place to excited state results in large change in energy.

Vibrational energy can be shown as:

E = hf (v)

Where, h is planks constant

f is frequency

v is vibrational energy

Step 2

As more rotation and vibration is taking place in a molecule than atom thus more electron transition is taking place in it. This results in appearance of more number of peaks and the rotational vibrational energy will get an additional factor of ½ to show increase in energy as follows:

E = hf (v + 1/2)

Step 3

The energy of vibration or hf can be shown graphically from transition l = 0 to l =1 as follows:

dl =-1 dl = +1 hf 1-1 to O 0 1:0 to 1

After going through the transition 0 to 1, the absorption of energy takes place by electron shown by right line in the graph and energy can be shown as:

E = hf + (l+1) h^2 /I

Where, l is angular momentum = 0,1 ,2